Convert 8.4 to a percent.
Answer by stanbon(57282)

Convert 8.4 to a percent.
-------
Percent means per hundred.
So write 8.4 with 100 in the denominator, as follows:
8.4/1 = [8.4*100]/[1*100] = 840/100 = 840%
=======================
Cheers,
Stan H.